OpenStudy (anonymous):

Find the total area of a cylinder whose radius is 2 cm and height is 4 cm

OpenStudy (anonymous):

Do you mean surface area?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

can u help me with it

OpenStudy (anonymous):

Okay. So, if the radius is 2, then the area of two circles is 2A= 2(pi*radius^2) 2A = 2(pi*4) 2A = 8pi and then there's the other part of the cylinder -> it's the circumference times the height so C = 2r*pi C = 2*2*pi C=4pi 4pi*4 = 16pi so 16pi+8pi = 24pi

OpenStudy (anonymous):

thats right thanks

OpenStudy (anonymous):

yw :)